# Investigating the Impact of Project Risks on Employee Turnover Intentions in the IT Industry of Pakistan

### v1

@@ 3.5 Organizational Context/ Macro Level
Researchers have also examined the impact of organizational diversity on employee turnover. Elvira and Cohen investigated the effect of gender composition on one’s own sex at various positions within the organization. Their results indicated that employees of the same sex had an impact on the retention of female employees, negatively affecting the turnover rate, but had no impact on male employees. Research on public accountants conducted by Saroj Parasuraman (1992) reported that female public accountants are more likely to leave their positions compared to their male counterparts, not due to family issues but because of the lack of promotional opportunities. Conversely, research results by Lyness (2001) showed that female financial managers were less likely to leave compared to male financial managers when they obtained a higher position in the preceding eleven months. Demographic misfit within work groups is also associated with a high turnover hazard. A comprehensive study conducted by Mallol (2007) on more than 20 US corporations and 450,000 professionals and managers found that incumbents of jobs typically held by African Americans and Hispanics were at a higher risk of turnover. Other organizational variables, such as firms’ pay distribution, also affect employee turnover Matt and John (2002)Eisenberger R1 (2002) found that the negative effect of perceived supervisory support on employee turnover is mediated by perceived organizational support. However, a study by 753 (a) showed that perceived supervisor support has an independent effect on turnover cognition, and perceived organizational support had a significant impact on turnover mediated through affective and organizational commitment.
@@ 5.5 Data Collected
Gender
Total Employees
Male
101
Female
49
@@
Age
Total Employees
Less than 25
62
25-29
61
30-39
21
>39
6
@@
Work Experience
Total Employees
< 5 years
73
5-10 years
44
10-15 years
22
16-20 years
10
@@
Salary
Total Employees
< 25000
34
25000-50000
61
51000-75000
26
76000-100000
17
>100000
12
@@ 8 Data reliability
Factor
Cronbach alpha score
Mean
Variance
Standard deviations
Questions
Delay in recruiting and resourcing
0.796
31.9265
30.875
5.55
9
Dependability risk
0.753
10.63
5.012
2.2387
3
Team composition risk
0.756
22.418
16.634
4.07
6
Control Process Risk
0.766
14.4615
9.627
3.10281
4
Job satisfaction
0.845
11.7353
6.824
2.61236
3
Perceived Job Alternatives
0.705
6.9918
3.276
1.80988
2
External Network
0.908
18.1765
25.730
5.07243
6
Internal Network
0.896
21.2059
22.255
4.717
6
Turnover intention
0.808
7.4853
6.880
2.6235
3
@@ 9.1 Correlation Analysis
Factors
Job Satisfaction
Team composition Risk
0.622*
Risk Evolved in estimating software requirements
0.578*
Control process risk
0.38
Dependability risk
0.501*
Perceived Job Alternatives
0.488*
Internal Network
0.548*
@@
Factors
Turnover Intentions
Job satisfaction
-0.560*
Perceived Job Alternatives
0.16
@@
Factors
Perceived Job Alternatives
Internal networking
0.200
External networking
0.24
@@ 9.2 Direct Regression Analysis
Variables
Job satisfaction
Column No
1 Beta
2 Beta
3 Beta
4 Beta
5 Beta
Age
0.14
0.09
0.07
0.15
0.06
Gender
0.07
0.1
0.17
0.09
0.10
Salary Range: <25000
0.18
0.12
0.11
0.19
0.16
25000 – 50000
0.21
0.18
0.13
0.15
0.19
51000 - 75000
0.23
0.11
0.16
0.21
0.20
76000 - 100000
0.27
0.28
0.222
0.23
0.18
>100000
0.34
0.30
0.33
0.27
0.26
Work Experience <5 years
0.16
0.13
0.21
0.18
0.17
5-10 years
0.22
0.21
0.14
0.21
0.18
10-15 years
0.29
0.27
0.11
0.17
0.23
16-20 years
0.31
0.30
0.16
0.19
0.21
TCR
0.622*
CPR
0.46
DR
0.499*
RIESR
0.578*
R square
0.481
0.073
0.362
0.584
R square change
0.472
0.087
0.359
0.578
@@ 5 Beta
Variables
Turnover Intentions
Column No
1 Beta
2 Beta
Age
0.10
0.13
Gender
0.07
0.09
Salary Range: <25000
0.18
0.19
25000 – 50000
0.20
0.21
51000 - 75000
0.22
0.23
76000 - 100000
0.27
0.25
>100000
0.29
0.26
Work Experience <5 years
0.11
0.17
5-10 years
0.12
0.21
10-15 years
0.14
0.22
16-20 years
0.13
0.27
JS
-0.560*
PJA
0.69*
R square
0.41
0.55
R square change
0.42
0.51
@@ 2 Beta
Variables
PJA
Column No
1 Beta
2 Beta
External Networking
0.16
Internal Networking
0.21
R square
0.018
0.026
R square change
0.014
0.027
@@ 9.3 Moderating Effect of perceived Job Alternatives:
Variables
STEP1
STEP2
STEP3
JS
-.565**
-0.578**
-0.618*
PJA
0.714**
0.456*
Interaction Term (JS * PJA)
0.095
R square
-.565
0.65
0.387
@@ 9.4 Meditation Test
Variables
Step1
Step2
Step3
Step4
CPR
-0.21
0.19
0.18
JS
-0.561**
-0.475**
R square
0.017
0.022
0.48
0.114
@@
Variables
Step1
Step2
Step3
Step4
DR
-0.501*
0.498**
-0.452**
JS
-0.582**
-0.356**
R square
0.369
0.378
0.49
0.46
@@
Variables
Step1
Step2
Step3
Step4
RIESR
-0.578*
0.547**
-0.571**
JS
-0.596**
-0.456**
R square
0.369
0.378
0.49
0.46
@@ 10 Conclusions
To gather data, we distributed a questionnaire to numerous companies across Pakistan, resulting in responses from 160 participants. Our analyses encompassed various statistical techniques, including regression analysis, step-wise linear regression, and correlation analysis, to evaluate our hypotheses. Our findings revealed that minimizing project-specific risks is essential for reducing turnover intentions. Specifically, all project-specific risk factors, with the exception of control process risk, mediated through job satisfaction and influenced employee turnover. Additionally, while perceived job alternatives did not moderate the relationship between job satisfaction and turnover intention, they independently influenced turnover intention. Similarly, our analyses indicated that external and internal networking did not directly impact turnover intention. In summary, this study sheds light on critical factors influencing turnover intentions in Pakistan’s IT industry and underscores the importance of managing project-specific risks and perceived job alternatives to mitigate turnover.

### v2

@@ 3.5 Organizational Context/ Macro Level
Researchers have also examined the impact of organizational diversity on employee turnover. Elvira and Cohen investigated the effect of gender composition on one’s own sex at various positions within the organization. Their results indicated that employees of the same sex had an impact on the retention of female employees, negatively affecting the turnover rate, but had no impact on male employees. Research on public accountants conducted by  Saroj Parasuraman (1992) reported that female public accountants are more likely to leave their positions compared to their male counterparts, not due to family issues but because of the lack of promotional opportunities. Conversely, research results by Lyness (2001) showed that female financial managers were less likely to leave compared to male financial managers when they obtained a higher position in the preceding eleven months. Demographic misfit within work groups is also associated with a high turnover hazard. A comprehensive study conducted by Mallol (2007) on more than 20 US corporations and 450,000 professionals and managers found that incumbents of jobs typically held by African Americans and Hispanics were at a higher risk of turnover. Other organizational variables, such as firms’ pay distribution, also affect employee turnover Matt and John (2002).  Eisenberger R1 (2002) found that the negative effect of perceived supervisory support on employee turnover is mediated by perceived organizational support. However, a study by  753 (a) showed that perceived supervisor support has an independent effect on turnover cognition, and perceived organizational support had a significant impact on turnover mediated through affective and organizational commitment.
@@ 5.5 Data Collected
Gender Total Employees
Male 101
Female 49
@@
Age Total Employees
Less than 25 62
25-29 61
30-39 21
>39 6
@@
Work Experience Total Employees
< 5 years 73
5-10 years 44
10-15 years 22
16-20 years 10
@@
Salary Total Employees
< 25000 34
25000-50000 61
51000-75000 26
76000-100000 17
>100000 12
@@ 8 Data reliability
Factor Cronbach alpha score Mean Variance Standard deviations Questions
Delay in recruiting and resourcing 0.79 31.92 30.87 5.55 9
Dependability risk 0.753 10.63 5.012 2.23 3
Team composition risk 0.756 22.41 16.63 4.07 6
Control Process Risk 0.76 14.46 9.62 3.10 4
Job satisfaction 0.84 11.73 6.82 2.61 3
Perceived Job Alternatives 0.70 6.99 3.27 1.80 2
External Network 0.90 18.17 25.73 5.072 6
Internal Network 0.89 21.20 22.25 4.717 6
Turnover intention 0.80 7.48 6.88 2.62 3
@@ 9.1 Correlation Analysis
Factors Job Satisfaction
Team composition Risk 0.622*
Risk Evolved in estimating software requirements 0.578*
Control process risk 0.38
Dependability risk 0.501*
Perceived Job Alternatives 0.488*
Internal Network 0.548*
@@
Factors Turnover Intentions
Job satisfaction -0.560*
Perceived Job Alternatives 0.16
@@
Factors Perceived Job Alternatives
Internal networking 0.200
External networking 0.24
@@ 9.2 Direct Regression Analysis
Variables Job satisfaction
Column No 1 Beta 2 Beta 3 Beta 4 Beta 5 Beta
Age 0.14 0.09 0.07 0.15 0.06
Gender 0.07 0.1 0.17 0.09 0.10
Salary Range: <25000 0.18 0.12 0.11 0.19 0.16
25000 – 50000 0.21 0.18 0.13 0.15 0.19
51000 - 75000 0.23 0.11 0.16 0.21 0.20
76000 - 100000 0.27 0.28 0.222 0.23 0.18
>100000 0.34 0.30 0.33 0.27 0.26
Work Experience <5 years 0.16 0.13 0.21 0.18 0.17
5-10 years 0.22 0.21 0.14 0.21 0.18
10-15 years 0.29 0.27 0.11 0.17 0.23
16-20 years 0.31 0.30 0.16 0.19 0.21
TCR - 0.622* - - -
CPR - - 0.46 - -
DR - - - 0.499* -
RIESR - - - - 0.578*
R square - 0.481 0.073 0.362 0.584
R square change - 0.472 0.087 0.359 0.578
@@
Variables Turnover Intentions
Column No 1 Beta 2 Beta
Age 0.10 0.13
Gender 0.07 0.09
Salary Range: <25000 0.18 0.19
25000 – 50000 0.20 0.21
51000 - 75000 0.22 0.23
76000 - 100000 0.27 0.25
>100000 0.29 0.26
Work Experience <5 years 0.11 0.17
5-10 years 0.12 0.21
10-15 years 0.14 0.22
16-20 years 0.13 0.27
JS -0.560* -
PJA - 0.69*
R square 0.41 0.55
R square change 0.42 0.51
@@
Variables PJA
Column No 1 Beta 2 Beta
External Networking 0.16 -
Internal Networking - 0.21
R square 0.018 0.026
R square change 0.014 0.027
@@ 9.3 Moderating Effect of perceived Job Alternatives:
Variables STEP1 STEP2 STEP3
JS -.565** -0.578** -0.618*
PJA - 0.714** 0.456*
Interaction Term (JS * PJA) 0.095
R square -.565 0.65 0.387
@@ 9.4 Meditation Test
Variables Step1 Step2 Step3 Step4
CPR -0.21 0.19 - 0.18
JS - - -0.561** -0.475**
R square 0.017 0.022 0.48 0.114
@@
Variables Step1 Step2 Step3 Step4
DR -0.501* 0.498** - -0.452**
JS - - -0.582** -0.356**
R square 0.369 0.378 0.49 0.46
@@
Variables Step1 Step2 Step3 Step4
RIESR -0.578* 0.547** - -0.571**
JS - - -0.596** -0.456**
R square 0.369 0.378 0.49 0.46
@@ 10 Conclusions
To gather data, we distributed a questionnaire to numerous companies across Pakistan, resulting in responses from 150 participants. Our analyses encompassed various statistical techniques, including regression analysis, step-wise linear regression, and correlation analysis, to evaluate our hypotheses. Our findings revealed that minimizing project-specific risks is essential for reducing turnover intentions. Specifically, all project-specific risk factors, with the exception of control process risk, mediated through job satisfaction and influenced employee turnover. Additionally, while perceived job alternatives did not moderate the relationship between job satisfaction and turnover intention, they independently influenced turnover intention. Similarly, our analyses indicated that external and internal networking did not directly impact turnover intention. In summary, this study sheds light on critical factors influencing turnover intentions in Pakistan’s IT industry and underscores the importance of managing project-specific risks and perceived job alternatives to mitigate turnover.